You are engineering a shipping operation for a major online retailer. Packages move on a conveyor belt. You have pulled 7 packages from the conveyor belt to weigh

The weights in pounds are as follows:

Weigh Pounds

3.4

3.1

2.0

4.3

2.9

4.5

4.1

Is weight a continuous variable or attribute? _____ (variable or attribute)

What is the sample average (x-bar)? _____ (one decimal will work)

What is the range? _____ (one decimal will work)

If packages above 4 pounds are "too heavy", what proportion of these packages is "too heavy?" _____ (two decimals)

Structured with the "too heavy" comment and if you desire to measure the packages as "too heavy" or "okay", is this now a variable or attribute? _____ (variable or attribute)

Do you really have enough data to do much of an analysis? (yes or no) ______
Answer #1

1.Is weight a continuous variable or attribute? _____ (variable or attribute)

Answer -Variable

Variables are usually much more dynamic and are usually used for doing calculations.

  • Anything that can be measured on a continuous basis.
  • Can always be divided into smaller increments.
  • Exists on a continuum
  • Preferred over Discrete.

2. What is the sample average (x-bar)?

=(3.4+3.1+2.0+4.3+2.9+4.5+4.1)/7

=3.5

3. What is the range?

Range = Maximum Value–Minimum Value

=4.5-2.0

=2.5

4.If packages above 4 pounds are "too heavy", what proportion of these packages is "too heavy?" _____ (two decimals)

Answer-Total number of packages=7

Number of packages above 4=3

Proportion = 3/7

                  = 0.42

                  = 42%

5.Structured with the "too heavy" comment and if you desire to measure the packages as "too heavy" or "okay", is this now a variable or attribute? _____ (variable or attribute)   

Answer- Variable

Variables are usually much more dynamic and are usually used for doing calculations:

6.Do you really have enough data to do much of an analysis? (yes or no) ______

Answer=yes enough data.

We can calculate mean,mode, median,range, Weighted mean.,Harmonic mean.,Geometric mean.,Root-Mean Square mean.